OpenStudy (errrwut):

Really need help** PDE Consider the following two-point boundary value problem in which  is a positive number. (equations inside) Use the Rayleigh Quotient to show that all eigenvalues are non negative. Explain why the number zero is not an eigenvalue. Find 2 × 2 matrices  and  so that conditions (ii) and (iii) are equivalent to (equations inside)

OpenStudy (errrwut):

(i) \[-\phi’’(x) = \lambda \phi(x),for -L \le x \le L\] (ii) \[\phi’(-L) = 0\] (iii) \[\phi(L) = 0.\] \[M\left(\begin{matrix}\phi(-L) \\ \phi'(-L)\end{matrix}\right) + N\left(\begin{matrix}\phi(L) \\ \phi’(L)\end{matrix}\right) = \left(\begin{matrix}0 \\ 0\end{matrix}\right)\]